1. A restaurant faces very high demand for its signature mousse desserts in the evening but is less busy during the day. Its manager estimates that inverse demand functions are pe = 34 - Qe in the evening and pd = 24 - Qd during the day, where e and d denote evening and daytime. The marginal cost of producing its dessert evening, MCe, is $8. The marginal cost of producing its dessert daytime, MCd, is $4. There is no fixed cost of producing dessert.

Create a spreadsheet with the column headings Qe, Pe, TRe, MRe, TCe, MCe, πe, Qd, Pd, TRd, MRd, TCd, MCd, and πd. (note: πe is profit evening and πd indicates profit daytime)

What is the optimal prices for the dessert that the restaurant should charge during the evening hours? __________

What is the optimal quantity for the dessert that the restaurant should produce during the evening hours? __________

What is the total cost of producing the optimal quantity for the dessert during the evening hours? __________

What is the maximum profit for the dessert that the restaurant should produce during the evening hours? __________

What is the optimal prices for the dessert that the restaurant should charge during the daytime hours? __________

What is the optimal quantity for the dessert that the restaurant should produce during the daytime hours? __________

What is the total cost of producing the optimal quantity for the dessert during the daytime hours? __________

What is the maximum profit for the dessert that the restaurant should produce during the daytime hours? __________

QUESTION 2

  1. A monopoly sells its good in the U.S. and Japanese markets. The American inverse demand function is PA =60−QA, and the Japanese inverse demand function is PJ =80−2QJ, where both prices, PA and PJ, are measured in dollars. The firm’s total cost of production is TC=5+16Q in both countries. Assume that the firm can prevent resale in other countries.

    What price will it charge in the U.S.? ________

    What is the optimal quantity in the U.S.? ________

    What is the total cost producing the optimal quantity in the U.S.? ________

    What is the maximum profit for the good in the U.S.? ________

    What price will it charge in the Japanese markets? ________

    What is the optimal quantity in the Japanese markets? ________

    What is the total cost producing the optimal quantity in the Japanese markets? ________

    What is the maximum profit for the good in the Japanese markets? ________

    Now, resale is allowed in both markets. What is the optimal price in both markets (round to the nearest whole number? ________

    What is the optimal quantity in both markets? ________
Answer #1
Qe Pe TRe MRe TCe MCe πe Qd Pd TRd MRd TCd MCd πd
0 34 0 0 0 0 24 0 0 0
1 33 33 33 8 8 25 1 23 23 23 4 4 19
2 32 64 31 16 8 48 2 22 44 21 8 4 36
3 31 93 29 24 8 69 3 21 63 19 12 4 51
4 30 120 27 32 8 88 4 20 80 17 16 4 64
5 29 145 25 40 8 105 5 19 95 15 20 4 75
6 28 168 23 48 8 120 6 18 108 13 24 4 84
7 27 189 21 56 8 133 7 17 119 11 28 4 91
8 26 208 19 64 8 144 8 16 128 9 32 4 96
9 25 225 17 72 8 153 9 15 135 7 36 4 99
10 24 240 15 80 8 160 10 14 140 5 40 4 100
11 23 253 13 88 8 165 11 13 143 3 44 4 99
12 22 264 11 96 8 168 12 12 144 1 48 4 96
13 21 273 9 104 8 169 13 11 143 -1 52 4 91
14 20 280 7 112 8 168 14 10 140 -3 56 4 84
15 19 285 5 120 8 165 15 9 135 -5 60 4 75
16 18 288 3 128 8 160 16 8 128 -7 64 4 64
17 17 289 1 136 8 153 17 7 119 -9 68 4 51
18 16 288 -1 144 8 144 18 6 108 -11 72 4 36
19 15 285 -3 152 8 133 19 5 95 -13 76 4 19
20 14 280 -5 160 8 120 20 4 80 -15 80 4 0
21 13 273 -7 168 8 105 21 3 63 -17 84 4 -21
22 12 264 -9 176 8 88 22 2 44 -19 88 4 -44
23 11 253 -11 184 8 69 23 1 23 -21 92 4 -69
24 10 240 -13 192 8 48 24 0 0 -23 96 4 -96
25 9 225 -15 200 8 25 25 -1 -25 -25 100 4 -125
26 8 208 -17 208 8 0 26 -2 -52 -27 104 4 -156
27 7 189 -19 216 8 -27 27 -3 -81 -29 108 4 -189
28 6 168 -21 224 8 -56 28 -4 -112 -31 112 4 -224
29 5 145 -23 232 8 -87 29 -5 -145 -33 116 4 -261
30 4 120 -25 240 8 -120 30 -6 -180 -35 120 4 -300
31 3 93 -27 248 8 -155 31 -7 -217 -37 124 4 -341
32 2 64 -29 256 8 -192 32 -8 -256 -39 128 4 -384
33 1 33 -31 264 8 -231 33 -9 -297 -41 132 4 -429
34 0 0 -33 272 8 -272 34 -10 -340 -43 136 4 -476

TRe = Pe*Qe = (34 – Qe) *Qe = 34Qe – Qe^2

MRe = dTRe/dQe = 34 – 2Qe

At equilibrium:

MRe = MCe:

34 – 2Qe = 8

2Qe = 26

Qe = 26/2 = 13.

Optimal Price from Demand curve:

Pe* = 34 – 13 = 21

Optimal quantity Qe* = 13.

Total cost of producing this quantity TCe* = 8*Qe*= 8*13 = 104.

Maximum profit = TRe – TCe = 13*21 – 8*13 =169.

During Day time:

MRd = dTRd/dQd

TRd = (24 – Qd)*Qd = 24Qd – Qd^2

MRd = 24 – 2Qd

Equilibrium:

MRd = MCd

24 – 2Qd = 4

2Qd = 20

Qd* = 10.

Optimal price from demand curve:

Pd* = 24 – 10 = 14

Optimal quantity: Qd* = 10.

Total cost = MC*Qd* = 4*10 = 40

Profit πd = TRe – TCe = 14*10 – 40 = 100

Hi. Please post questions separately.